OpenStudy (anonymous):

the the area of a rectangle is 16 square feet. determine the length and width if the length is 4times the width.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 4x(x) = 16 }\) \(\Large \color{purple}{\rightarrow 4x^2 = 16 }\) \(\Large \color{purple}{\rightarrow x^2 = 4 }\) Get the sqrt of both sides(easy lol)

Parth (parthkohli):

Concept insight \(\Large \color{purple}{\rightarrow Area(Rectangle) = l*b }\)

OpenStudy (anonymous):

length is 16 width is 4

Parth (parthkohli):

No, no, no. x is the width. x = 2 and length(4x) = 8

OpenStudy (anonymous):

okay

Parth (parthkohli):

We can confirm by multiplying 8 and 2. We get 16 square feet :D

OpenStudy (anonymous):

kay- thanks

Parth (parthkohli):

np